In a university class that includes 120 students, mostly non-traditional, the Mean age is 32, with a standard deviation 12 students, and with a normal distribution. Six of these students are international students What is the probability that, if one student is picked at a random, he/she be a teenager (12 to 19 years old)? You need to (somewhat) critically think about this before tackling it!! 1. What is the probability that if ONE student is selected at random, he/she will be an international student? 2. 3. What is the probability that if ONE student is selected at random, his/her age be between 28 to 30 years old? What is the probability that if on student is selected at random, his/her age be between 28 to 36 years old? 4. 5. Suppose we select, at random, a sample of 9 students from this class. What is the probability that the MEAN age of that sample will be more than 38 years old?
Answer #1

Solution :-

Given data :-

Number of students = 120

Mean age, mu = 32

standard deviation, s = 12

( 1 ) :-

Here we need to find out the probability that, if one student is picked at a random, he/she be a teenagers ( 12 to 19 years old).

i.e, P(12 < X < 19)

P(12< X < 19)P(X < 19)- P(X < 12)

P(12 < X < 19)-P

19 32 12- 32 P(12 X 19) P

P(12leq Xleq 19)=Pleft (rac{13}{12} ight )-Pleft (rac{-20 }{12} ight )

P(12leq Xleq 19)=Pleft (-1.083 ight )-Pleft (1.667 ight )

P(12 < X < 19)-0.14007-0.04846 ( ecause From standard normal table)

P(12leq Xleq 19)=0.09161      

( 2 ) :-

Here we need to find out the probability that if ONE student is selected at random, he/she will be an international student.

i.e, P(International student )

P(International student ) = 6/120

P(International student ) = 0.05

( 3 ) :-

Here we need to find out the probability that if on student is selected at random, his/her age be between 28 to 30 years old.

i.e, P(28leq Xleq 30)

P(28leq Xleq 30)=Pleft (z<rac{30-mu}{s} ight ) -Pleft ( z<rac{28-mu}{s} ight )

P(28leq Xleq 30)=Pleft (z<rac{30-32}{12} ight ) -Pleft ( z<rac{28-32}{12} ight )

P(28leq Xleq 30)=Pleft (z<rac{-2}{12} ight ) -Pleft (z< rac{-4}{12} ight )

P(28leq Xleq 30)=Pleft (z<-0.1667 ight ) -Pleft ( z<-0.333 ight )

P(28leq Xleq 30)=0.43644 -0.37070    ( ecause From standard normal table)

P(28leq Xleq 30)=0.06574

( 4 ) :-

Here we need to find out the probability that if on student is selected at random, his/her age be between 28 to 36 years old.

i.e, P(28 < X < 36)

P(28leq Xleq 36)=P(Xleq 36)-P(Xleq 28)

P(28leq Xleq 36)=Pleft (z<rac{36-mu }{s} ight )-Pleft (z<rac{28-mu }{s} ight )

P(28leq Xleq 36)=Pleft (z<rac{36-32 }{12} ight )-Pleft (z<rac{28-32 }{12} ight )

P(12leq Xleq 19)=Pleft (z<rac{4}{12} ight )-Pleft (z<rac{-4 }{12} ight )

P(12leq Xleq 19)=Pleft (z<0.334 ight )-Pleft (z<-0.334 ight )

P(12leq Xleq 19)=0.62930 -0.37070    ( ecause From standard normal table)

P(12leq Xleq 19)=0.2686

( 5 ) :-

Here we need to find out the probability that the mean age of that sample will be more than 38 years old.

i.e, P(xgeq 38)

P(xgeq 38)=Pleft (z>rac{38-mu }{s/sqrt{n}} ight )

P(xgeq 38)=Pleft (z>rac{38-32 }{12/sqrt{9}} ight )

P(xgeq 38)=Pleft (z>rac{6 }{4} ight )

P(xgeq 38)=Pleft (z>1.5 ight )

P(xgeq 38)=1-Pleft (z<1.5 ight )

P(xgeq 38)=1-0.9332    ( ecause From standard normal table)

P(xgeq 38)=0.0668​​​​​​​